[figure: 9.7 cm]

ELECTRONIC STRUCTURE OF THE URANIUM MONOSTANNIDE USn

A. Szajek Institute of Molecular Physics, Polish Academy of Sciences, Poznań Poland

The uranium monostannide USn crystallizes in the orthorhombic T hIn-type structure with the space group P bcm [1]. Uranium and tin atoms each occupy three crystallographic positions. The uranium atoms located in (4c) sites form metallic chains running along c axis, with a U − U interatomic distance of 3.10 Å. USn orders ferromagnetically below T_{C} = 55K, with an ordered moment of 1.20μ_{B}/U measured on a polycrystallized sample. Because of short distance U(4c) − U(4c) and expected hybridization with the conduction-band electron states it would be interesting to know if this uranium sublattice is also ordered magnetically. Ab − initio calculations would clarify this situation. In this work, spin-polarized band structure calculations results will be presented based on the tight binding version of the linear muffin-tin orbital method in the atomic sphere approximation (TB LMTO ASA). The calculations showed that all three types of uranium atoms are magnetically ordered and their moments depend on local environments. Work supported by the KBN Grant No. 2 P03B 023 22

[1] P. Boulet and H. Noël, Solid State Commun. 107 (1998) 135

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : poster

Corresponding author : A. Szajek

Address for correspondence : Institute of Molecular Physics Polish Academy of Sciences ul. Smoluchowskiego 17, 60-179 Poznań, Poland

Email address : szajek@ifmpan.poznan.pl